# N. Deng & Y. Wang & W. Cao

### v1

@@ 3.1 Key estimates for two auxiliary equations
$\leq$ $-\int_{\mathcal{O}}y^{2}(t,x)|\partial_{x}y(t,x)|^{2}\mathrm{d}x+C\int_{\mathcal{O}}\left(|z(t,x)\partial_{x}z(t,x)y(t,x)\partial_{x}y(t,x)|+|z^{2}(t,x)y(t,x)\partial_{xx}y(t,x)|+y^{2}(t,x)|\partial_{x}z(t,x)|^{2}\right)\mathrm{d}x$
@@
Substituting the above estimations into (20) yields
@@
$\|Y_{t}\|^{2}_{L^{2}}+\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\|AY_{s}\|^{2}_{L^{2}}\mathrm{d}s$ (21)
@@
We next estimate $\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left\|Y_{s}+Z_{s}\right\|^{4}_{L^{4}}\mathrm{d}s$. According to (17), we have $\left\langle Y_{t},\phi_{0}\right\rangle_{L^{2}}=\left\langle y_{0},\phi_{0}\right\rangle_{L^{2}}$ and
@@
$\sum_{j=1}^{\infty}\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\left\langle Y_{t}-y_{0},\phi_{j}\right\rangle^{2}_{L^{2}}}{2\lambda_{j}}\right)+\big\|(-A)^{\frac{1}{2}}Y_{t}\big\|^{2}_{L^{2}}$ (22)
@@
$\big\langle F(Y_{t}+Z_{t}),Y_{t}+Z_{t}\big\rangle\geq\frac{7}{8}a_{0}\left\|Y_{t}+Z_{t}\right\|_{L^{4}}^{4}-C,$ (23)
@@
$\big\langle F(Y_{t}+Z_{t}),Z_{t}+y_{0}\big\rangle\leq$ $\frac{5}{8}a_{0}\left\|Y_{t}+Z_{t}\right\|_{L^{4}}^{4}+C\left(1+\|y_{0}\|_{L^{\infty}}^{4}+\left\|Z_{t}\right\|_{L^{4}}^{4}\right).$ (24)
Substituting (23) and (24) into (22), then applying Hölder’s inequality and Young’s inequality, we obtain
$\sum_{j=1}^{\infty}\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\left\langle Y_{t}-y_{0},\phi_{j}\right\rangle^{2}_{L^{2}}}{2\lambda_{j}}\right)\leq$ $-\frac{1}{2}\big\|(-A)^{\frac{1}{2}}Y_{t}\big\|^{2}_{L^{2}}-\frac{a_{0}}{4}\left\|Y_{t}+Z_{t}\right\|_{L^{4}}^{4}+C\left(1+\|y_{0}\|_{L^{\infty}}^{4}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}^{2}+\left\|Z_{t}\right\|_{L^{4}}^{4}\right).$ (25)
@@
This combined with (21) verifies (19).
@@
$\big\|(-A)^{\frac{1}{2}}Y_{t}\big\|_{l^{2}_{N}}\leq$ $C\left(1+\|y_{0}\|_{L^{\infty}}^{20}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}^{10}+\mathbb{I}_{1}(t,3/4,\lambda_{1}^{2}/2,|K_{1}|^{3}+\left\|Z\right\|_{L^{6}}^{3})+\mathbb{I}_{2}(t,3/4,\lambda_{1}^{2}/2,|K_{1}|^{5}+\left\|Z\right\|_{L^{6}}^{9})\right).$ (26)
The mild solution of (17) is given by
@@
Using (4), we have
$\big\|(-A)^{\frac{1}{2}}Y_{t}\big\|_{L^{2}}\leq$ $\big\|(-A)^{\frac{1}{2}}e^{-A^{2}t}y_{0}\big\|_{L^{2}}+\int_{0}^{t}\big\|(-A)^{\frac{3}{2}}e^{-A^{2}(t-s)}F(Y_{s}+Z_{s})\big\|_{L^{2}}\mathrm{d}s$ (27)
@@
Applying (15), (16), (4), and Hölder’s inequality, one can obtain
@@
$\leq$ $\big\|e^{-A^{2}t}y_{0}\big\|_{L^{\infty}}+C\int_{0}^{t}\left(\big\|A^{2}e^{-A^{2}(t-s)}F(Y_{s}+Z_{s})\big\|_{L^{2}}^{\frac{1}{6}}\big\|Ae^{-A^{2}(t-s)}F(Y_{s}+Z_{s})\big\|_{L^{2}}^{\frac{5}{6}}+\big\|Ae^{-A^{2}(t-s)}F(Y_{s}+Z_{s})\big\|_{L^{2}}\right)\mathrm{d}s$
$\leq$ $C\bigg(\|y_{0}\|_{L^{2}}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}+\int_{0}^{t}\left((t-s)^{-\frac{7}{12}}+(t-s)^{-\frac{1}{2}}\right)e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left\|F(Y_{s}+Z_{s})\right\|_{L^{2}}\mathrm{d}s\bigg).$ (28)
By (15) and Young’s inequality, we derive that for any $0\leq\alpha<\frac{3}{4}$,
$\int_{0}^{t}(t-s)^{-\alpha}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left\|F(Y_{s}+Z_{s})\right\|_{L^{2}}\mathrm{d}s$ (29)
@@
By combining (28), (29), and Hölder’s inequality, we obtain
$\left\|Y_{t}\right\|_{L^{6}}^{3}\leq$ $C\bigg(1+\|y_{0}\|_{L^{2}}^{3}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}^{3}\bigg)+C\left|\int_{0}^{t}\left((t-s)^{-\frac{7}{12}}+(t-s)^{-\frac{1}{2}}\right)e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left(\left\|Y_{s}\right\|_{L^{2}}^{3}+\left\|Z_{s}\right\|_{L^{2}}^{3}\right)\mathrm{d}s\right|^{3}$ (30)
@@
It follows from (19), (27), and (30) that (26) holds.
Finally, by gathering (19) and (26), we obtain (18). This completes the proof.
@@
A similar estimate for the perturb equation in the $\dot{H}^{-1}$-norm can be found in Da Prato & Debussche (1996, Theorem 3.1). Nevertheless, such an estimate is insufficient to guarantee the existence of invariant measure on the space $H_{\alpha}$. In Proposition 3.2, we improve this estimate to establish the uniform moment boundedness of $u(t,\cdot)$ in the more stringent $L^{\infty}$-norm.
We next consider another auxiliary equation, which is used to prove the Feller property of $u(t,x)$. For any $\hat{u}_{0},\ \tilde{u}_{0}\in H_{\alpha}$, we let $U_{\hat{u}_{0},t}:=u(\hat{u}_{0};t,\cdot)$ and $U_{\tilde{u}_{0},t}:=u(\tilde{u}_{0};t,\cdot)$ denote the solutions of (1) with initial conditions $\hat{u}_{0}$ and $\tilde{u}_{0}$, respectively. Define $\hat{E}_{t}=U_{\hat{u}_{0},t}-U_{\tilde{u}_{0},t}$, which solves
$\left\{\begin{aligned} &\mathrm{d}\hat{E}_{t}+A^{2}\hat{E}_{t}\mathrm{d}t=A\big(F(\hat{E}_{t}+U_{\tilde{u}_{0},t})-F(U_{\tilde{u}_{0},t})\big)\mathrm{d}t,\quad t\in(0,\infty),\\ &\hat{E}_{0}=\hat{u}_{0}-\tilde{u}_{0}.\end{aligned}\right$. (31)
In the next proposition, we state a general result. In particular, when $\hat{X}_{t}=0$, $\hat{Z}_{t}=U_{\tilde{u}_{0},t}$, and $\hat{y}_{0}=\hat{u}_{0}-\tilde{u}_{0}$ in (32), we obtain the estimate for (31).
@@
Suppose that Assumption 2 holds and $\hat{Y}_{t}$ solves
$\left\{\begin{aligned} &\mathrm{d}\hat{Y}_{t}+A^{2}\hat{Y}_{t}\mathrm{d}t=A\left(F(\hat{Y}_{t}+\hat{X}_{t}+\hat{Z}_{t})-F(\hat{X}_{t})\right)\mathrm{d}t,\quad t\in(0,\infty),\\ &\hat{Y}_{0}=\hat{y}_{0},\end{aligned}\right$. (32)
@@
$\|\hat{Y}_{t}\|_{L^{p}(\Omega;H)}$ $\leq Ce^{-\frac{\epsilon_{0}}{2}t}\,\|\hat{y}_{0}\|_{L^{2}}+C\sqrt{K_{3}(t)}\,\Big(\mathbb{I}_{1}\big(t,1/2,1/2,\|\hat{Z}\|_{L^{2p}(\Omega;L^{2}(\mathcal{O}))}^{2}\big)\Big)^{\frac{1}{2}}$ (33)
@@
where $\epsilon_{0}$ is given in (8).
Proof 3.6.
It follows from (32) that for $t>0$, $\left\langle\hat{Y}_{t},\phi_{0}\right\rangle_{L^{2}}=0$ and
@@
It follows from Assumption 2, Young’s inequality, (7) and Hölder’s inequality that
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\sum_{j=1}^{\infty}\frac{\big\langle\hat{Y}_{t},\phi_{j}\big\rangle^{2}_{L^{2}}}{2\lambda_{j}}\right)=$ $-\big\|(-A)^{\frac{1}{2}}\hat{Y}_{t}\big\|^{2}_{L^{2}}+\Big\langle F(\hat{X}_{t})-F(\hat{X}_{t}+\hat{Y}_{t}),\hat{Y}_{t}\Big\rangle_{L^{2}}+\Big\langle F(\hat{X}_{t}+\hat{Y}_{t})-F(\hat{X}_{t}+\hat{Y}_{t}+\hat{Z}_{t}),\hat{Y}_{t}\Big\rangle_{L^{2}}$
@@
$\left\|\int_{0}^{t}e^{-\epsilon_{0}(t-s)}\big\|(-A)^{\frac{1}{2}}\hat{Y}_{s}\big\|^{2}_{L^{2}}\mathrm{d}s\right\|_{L^{p}(\Omega)}\leq$ $Ce^{-\epsilon_{0}t}\|\hat{y}_{0}\|_{L^{2}}^{2}+CK_{2}(t)\left(\int_{0}^{t}e^{-\epsilon_{0}(t-s)}\big\|\hat{Z}_{s}\big\|_{L^{2p}(\Omega;L^{4}(\mathcal{O}))}^{4}\mathrm{d}s\right)^{\frac{1}{2}},$ (34)
@@
$\|(-A)^{\frac{1}{2}}(v_{1}v_{2})\|_{L^{2}}\leq C\|v_{1}\|_{W^{1,2}}\|(-A)^{\frac{1}{2}}v_{2}\|_{L^{2}},\quad v_{1},v_{2}\in W^{1,2}(\mathcal{O})\ \text{and}\ \langle v_{2},\phi_{0}\rangle_{L^{2}}=0.$ (35)
Since the mild solution to (32) can be written as
@@
it can be derived from (4), (7) and (35) that
$\big\|\hat{Y}_{t}\big\|_{L^{2}}\leq$ $\big\|e^{-A^{2}t}\hat{y}_{0}\big\|_{L^{2}}+\int_{0}^{t}\big\|Ae^{-A^{2}(t-s)}\big(F(\hat{X}_{t}+\hat{Y}_{t}+\hat{Z}_{t})-F(\hat{X}_{t})\big)\big\|_{L^{2}}\mathrm{d}s$ (36)
@@
By applying the Hölder inequality, (36), (34), and using the Sobolev embedding $W^{1,2}(\mathcal{O})\hookrightarrow L^{\infty}(\mathcal{O})$, we obtain
@@
$\leq$ $Ce^{-\frac{\lambda_{1}^{2}}{2}t}\left\|\hat{y}_{0}\right\|_{L^{2}}+C\int_{0}^{t}(t-s)^{-\frac{1}{2}}e^{-\frac{\lambda_{1}^{2}}{2}t}\left(1+\big\|\hat{X}_{s}+\hat{Y}_{s}+\hat{Z}_{s}\big\|_{L^{2p}(\Omega;L^{\infty}(\mathcal{O}))}^{2}+\big\|\hat{X}_{s}+\hat{Y}_{s}\big\|_{L^{2p}(\Omega;L^{\infty}(\mathcal{O}))}^{2}\right)\big\|\hat{Z}_{s}\big\|_{L^{2p}(\Omega;H)}\mathrm{d}s$
$+C\left(\int_{0}^{t}(t-s)^{-\frac{1}{2}}e^{-(\lambda_{1}^{2}-\epsilon_{0})(t-s)}\left(1+\|\hat{X}_{s}+\hat{Y}_{s}\|_{L^{p}(\Omega;W^{1,2}(\mathcal{O}))}^{4}+\|\hat{X}_{s}\|_{L^{p}(\Omega;W^{1,2}(\mathcal{O}))}^{4}\right)\mathrm{d}s\right)^{\frac{1}{2}}\left\|\int_{0}^{t}e^{-\epsilon_{0}(t-s)}\left\|(-A)^{\frac{1}{2}}\hat{Y}_{s}\right\|_{L^{2}}\mathrm{d}s\right\|_{L^{p}(\Omega)}^{\frac{1}{2}}$
$\leq$ $Ce^{-\frac{\epsilon_{0}}{2}t}\left\|\hat{y}_{0}\right\|_{L^{2}}+C\sqrt{K_{3}(t)}\left(\mathbb{I}_{1}\Big(t,1/2,1/2,\big\|\hat{Z}\big\|_{L^{2p}(\Omega;L^{2}(\mathcal{O}))}^{2}\Big)\right)^{\frac{1}{2}}+C\sqrt{K_{2}(t)K_{3}(t)}\left(\mathbb{I}_{1}\Big(t,0,\epsilon_{0},\big\|\hat{Z}\big\|_{L^{2p}(\Omega;L^{4}(\mathcal{O}))}^{4}\Big)\right)^{\frac{1}{4}}.$
This completes the proof.
@@
Suppose that Assumptions 2-2 hold, then the transition semigroup $\{P_{t}\}_{t\in[0,\infty)}$ of SCHE (1) possesses the Feller property.
Proof 3.8.
For any $t\in[0,\infty)$, it follows from (31) and Proposition 3.5 that
$\left\|u(\hat{u}_{0};t,\cdot)-u(\tilde{u}_{0};t,\cdot)\right\|_{L^{2}(\Omega;H)}\leq Ce^{-\frac{\epsilon_{0}}{2}t}\left\|\hat{u}_{0}-\tilde{u}_{0}\right\|_{L^{2}},\quad\hat{u}_{0},\tilde{u}_{0}\in H.$ (37)
@@
By using the continuity and boundedness of $\phi$, (37), and the Hölder inequality, we obtain
@@
where ${B}^{C}$ denotes the complement of a set B{B}, and $\chi_{{B}}$ denotes its indicator function. Therefore, there exists a constant $\hat{\delta}>0$, such that, whenever $\|\hat{u}_{0}-\tilde{u}_{0}\|_{L^{2}}<\hat{\delta}$, we have $|P_{t}\phi(u_{0})-P_{t}\phi(\hat{u}_{0})|<\epsilon,$ which means $P_{t}$ is continuous. It follows from the boundedness of $\phi$ that $P_{t}$ is bounded. In summary, $\{P_{t}\}_{t\in[0,\infty)}$ possesses the Feller property
3.2 Existence and uniqueness of invariant measure

### v3

@@ 3.  Invariant measure for the stochastic Cahn–Hilliard equation
$\leq$ $-\int_{\mathcal{O}}y^{2}(t,x)|\partial_{x}y(t,x)|^{2}\mathrm{d}x$
$+C\int_{\mathcal{O}}\left(|z(t,x)\partial_{x}z(t,x)y(t,x)\partial_{x}y(t,x)|+|z^{2}(t,x)y(t,x)\partial_{xx}y(t,x)|+y^{2}(t,x)|\partial_{x}z(t,x)|^{2}\right)\mathrm{d}x$
@@
Substituting the above estimations into (3.12) yields
@@
$\|Y_{t}\|^{2}_{L^{2}}+\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\|AY_{s}\|^{2}_{L^{2}}\mathrm{d}s$ (3.13)
@@
We next estimate $\int_{0}^{t}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left\|Y_{s}+Z_{s}\right\|^{4}_{L^{4}}\mathrm{d}s$. According to (3.9), we have $\left\langle Y_{t},\phi_{0}\right\rangle_{L^{2}}=\left\langle y_{0},\phi_{0}\right\rangle_{L^{2}}$ and
@@
$\sum_{j=1}^{\infty}\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\left\langle Y_{t}-y_{0},\phi_{j}\right\rangle^{2}_{L^{2}}}{2\lambda_{j}}\right)+\big\|(-A)^{\frac{1}{2}}Y_{t}\big\|^{2}_{L^{2}}$ (3.14)
@@
$\big\langle F(Y_{t}+Z_{t}),Y_{t}+Z_{t}\big\rangle\geq\frac{7}{8}a_{0}\left\|Y_{t}+Z_{t}\right\|_{L^{4}}^{4}-C,$ (3.15)
@@
$\big\langle F(Y_{t}+Z_{t}),Z_{t}+y_{0}\big\rangle\leq$ $\frac{5}{8}a_{0}\left\|Y_{t}+Z_{t}\right\|_{L^{4}}^{4}+C\left(1+\|y_{0}\|_{L^{\infty}}^{4}+\left\|Z_{t}\right\|_{L^{4}}^{4}\right).$ (3.16)
Substituting (3.15) and (3.16) into (3.14), then applying Hölder’s inequality and Young’s inequality, we obtain
$\sum_{j=1}^{\infty}\frac{\mathrm{d}}{\mathrm{d}t}\left(\frac{\left\langle Y_{t}-y_{0},\phi_{j}\right\rangle^{2}_{L^{2}}}{2\lambda_{j}}\right)\leq$ $-\frac{1}{2}\big\|(-A)^{\frac{1}{2}}Y_{t}\big\|^{2}_{L^{2}}-\frac{a_{0}}{4}\left\|Y_{t}+Z_{t}\right\|_{L^{4}}^{4}+C\left(1+\|y_{0}\|_{L^{\infty}}^{4}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}^{2}+\left\|Z_{t}\right\|_{L^{4}}^{4}\right).$ (3.17)
@@
This combined with (3.13) verifies (3.11).
@@
$\big\|(-A)^{\frac{1}{2}}Y_{t}\big\|_{l^{2}_{N}}\leq$ $C\Big(1+\|y_{0}\|_{L^{\infty}}^{20}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}^{10}+\mathbb{I}_{1}(t,3/4,\lambda_{1}^{2}/2,|K_{1}|^{3}+\left\|Z\right\|_{L^{6}}^{3})$ (3.18)
$\qquad\qquad+\mathbb{I}_{2}(t,3/4,\lambda_{1}^{2}/2,|K_{1}|^{5}+\left\|Z\right\|_{L^{6}}^{9})\Big).$
The mild solution of (3.9) is given by
@@
Using (2.3), we have
$\big\|(-A)^{\frac{1}{2}}Y_{t}\big\|_{L^{2}}\leq$ $\big\|(-A)^{\frac{1}{2}}e^{-A^{2}t}y_{0}\big\|_{L^{2}}+\int_{0}^{t}\big\|(-A)^{\frac{3}{2}}e^{-A^{2}(t-s)}F(Y_{s}+Z_{s})\big\|_{L^{2}}\mathrm{d}s$ (3.19)
@@
Applying (3.7), (3.8), (2.3), and Hölder’s inequality, one can obtain
@@
$\leq$ $\big\|e^{-A^{2}t}y_{0}\big\|_{L^{\infty}}+C\int_{0}^{t}\Big(\big\|A^{2}e^{-A^{2}(t-s)}F(Y_{s}+Z_{s})\big\|_{L^{2}}^{\frac{1}{6}}\big\|Ae^{-A^{2}(t-s)}F(Y_{s}+Z_{s})\big\|_{L^{2}}^{\frac{5}{6}}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\big\|Ae^{-A^{2}(t-s)}F(Y_{s}+Z_{s})\big\|_{L^{2}}\Big)\mathrm{d}s$
$\leq$ $C\bigg(\|y_{0}\|_{L^{2}}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}+\int_{0}^{t}\left((t-s)^{-\frac{7}{12}}+(t-s)^{-\frac{1}{2}}\right)e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left\|F(Y_{s}+Z_{s})\right\|_{L^{2}}\mathrm{d}s\bigg).$ (3.20)
By (3.7) and Young’s inequality, we derive that for any $0\leq\alpha<\frac{3}{4}$,
$\int_{0}^{t}(t-s)^{-\alpha}e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left\|F(Y_{s}+Z_{s})\right\|_{L^{2}}\mathrm{d}s$ (3.21)
@@
By combining (3.20), (3.21), and Hölder’s inequality, we obtain
$\left\|Y_{t}\right\|_{L^{6}}^{3}\leq$ $C\bigg(1+\|y_{0}\|_{L^{2}}^{3}+\big\|(-A)^{\frac{1}{2}}y_{0}\big\|_{L^{2}}^{3}\bigg)+C\left|\int_{0}^{t}\left((t-s)^{-\frac{7}{12}}+(t-s)^{-\frac{1}{2}}\right)e^{-\frac{\lambda_{1}^{2}}{2}(t-s)}\left(\left\|Y_{s}\right\|_{L^{2}}^{3}+\left\|Z_{s}\right\|_{L^{2}}^{3}\right)\mathrm{d}s\right|^{3}$ (3.22)
@@
It follows from (3.11), (3.19), and (3.22) that (3.18) holds.
Finally, by gathering (3.11) and (3.18), we obtain (3.10). This completes the proof. ∎
@@
A similar estimate for the perturb equation in the $\dot{H}^{-1}$-norm can be found in [15, Theorem 3.1]. Nevertheless, such an estimate is insufficient to guarantee the existence of invariant measure on the space $H_{\alpha}$. In Proposition 3.2, we improve this estimate to establish the uniform moment boundedness of $u(t,\cdot)$ in the more stringent $L^{\infty}$-norm.
We next consider another auxiliary equation, which is used to prove the Feller property of $u(t,x)$. For any $\hat{u}_{0},\ \tilde{u}_{0}\in H_{\alpha}$, we let $U_{\hat{u}_{0},t}:=u(\hat{u}_{0};t,\cdot)$ and $U_{\tilde{u}_{0},t}:=u(\tilde{u}_{0};t,\cdot)$ denote the solutions of (1.1) with initial conditions $\hat{u}_{0}$ and $\tilde{u}_{0}$, respectively. Define $\hat{E}_{t}=U_{\hat{u}_{0},t}-U_{\tilde{u}_{0},t}$, which solves
$\left\{\begin{aligned} &\mathrm{d}\hat{E}_{t}+A^{2}\hat{E}_{t}\mathrm{d}t=A\big(F(\hat{E}_{t}+U_{\tilde{u}_{0},t})-F(U_{\tilde{u}_{0},t})\big)\mathrm{d}t,\quad t\in(0,\infty),\\ &\hat{E}_{0}=\hat{u}_{0}-\tilde{u}_{0}.\end{aligned}\right$. (3.23)
In the next proposition, we state a general result. In particular, when $\hat{X}_{t}=0$, $\hat{Z}_{t}=U_{\tilde{u}_{0},t}$, and $\hat{y}_{0}=\hat{u}_{0}-\tilde{u}_{0}$ in (3.24), we obtain the estimate for (3.23).
@@
Suppose that Assumption 2.2 holds and $\hat{Y}_{t}$ solves
$\left\{\begin{aligned} &\mathrm{d}\hat{Y}_{t}+A^{2}\hat{Y}_{t}\mathrm{d}t=A\left(F(\hat{Y}_{t}+\hat{X}_{t}+\hat{Z}_{t})-F(\hat{X}_{t})\right)\mathrm{d}t,\quad t\in(0,\infty),\\ &\hat{Y}_{0}=\hat{y}_{0},\end{aligned}\right$. (3.24)
@@
$\|\hat{Y}_{t}\|_{L^{p}(\Omega;H)}$ $\leq Ce^{-\frac{\epsilon_{0}}{2}t}\,\|\hat{y}_{0}\|_{L^{2}}+C\sqrt{K_{3}(t)}\,\Big(\mathbb{I}_{1}\big(t,1/2,1/2,\|\hat{Z}\|_{L^{2p}(\Omega;L^{2}(\mathcal{O}))}^{2}\big)\Big)^{\frac{1}{2}}$ (3.25)
@@
where $\epsilon_{0}$ is given in (2.7).
Proof.
It follows from (3.24) that for $t>0$, $\left\langle\hat{Y}_{t},\phi_{0}\right\rangle_{L^{2}}=0$ and
@@
It follows from Assumption 2.2, Young’s inequality, (2.6) and Hölder’s inequality that
$\frac{\mathrm{d}}{\mathrm{d}t}\left(\sum_{j=1}^{\infty}\frac{\big\langle\hat{Y}_{t},\phi_{j}\big\rangle^{2}_{L^{2}}}{2\lambda_{j}}\right)=$ $-\big\|(-A)^{\frac{1}{2}}\hat{Y}_{t}\big\|^{2}_{L^{2}}+\Big\langle F(\hat{X}_{t})-F(\hat{X}_{t}+\hat{Y}_{t}),\hat{Y}_{t}\Big\rangle_{L^{2}}$
$+\Big\langle F(\hat{X}_{t}+\hat{Y}_{t})-F(\hat{X}_{t}+\hat{Y}_{t}+\hat{Z}_{t}),\hat{Y}_{t}\Big\rangle_{L^{2}}$
@@
$\left\|\int_{0}^{t}e^{-\epsilon_{0}(t-s)}\big\|(-A)^{\frac{1}{2}}\hat{Y}_{s}\big\|^{2}_{L^{2}}\mathrm{d}s\right\|_{L^{p}(\Omega)}\leq$ $Ce^{-\epsilon_{0}t}\|\hat{y}_{0}\|_{L^{2}}^{2}+CK_{2}(t)\left(\int_{0}^{t}e^{-\epsilon_{0}(t-s)}\big\|\hat{Z}_{s}\big\|_{L^{2p}(\Omega;L^{4}(\mathcal{O}))}^{4}\mathrm{d}s\right)^{\frac{1}{2}},$ (3.26)
@@
$\|(-A)^{\frac{1}{2}}(v_{1}v_{2})\|_{L^{2}}\leq C\|v_{1}\|_{W^{1,2}}\|(-A)^{\frac{1}{2}}v_{2}\|_{L^{2}},\quad v_{1},v_{2}\in W^{1,2}(\mathcal{O})\ \text{and}\ \langle v_{2},\phi_{0}\rangle_{L^{2}}=0.$ (3.27)
Since the mild solution to (3.24) can be written as
@@
it can be derived from (2.3), (2.6) and (3.27) that
$\big\|\hat{Y}_{t}\big\|_{L^{2}}\leq$ $\big\|e^{-A^{2}t}\hat{y}_{0}\big\|_{L^{2}}+\int_{0}^{t}\big\|Ae^{-A^{2}(t-s)}\big(F(\hat{X}_{t}+\hat{Y}_{t}+\hat{Z}_{t})-F(\hat{X}_{t})\big)\big\|_{L^{2}}\mathrm{d}s$ (3.28)
@@
By applying the Hölder inequality, (3.28), (3.26), and using the Sobolev embedding $W^{1,2}(\mathcal{O})\hookrightarrow L^{\infty}(\mathcal{O})$, we obtain
@@
$\leq$ $Ce^{-\frac{\lambda_{1}^{2}}{2}t}\left\|\hat{y}_{0}\right\|_{L^{2}}+C\int_{0}^{t}(t-s)^{-\frac{1}{2}}e^{-\frac{\lambda_{1}^{2}}{2}t}\Big(1+\|\hat{X}_{s}+\hat{Y}_{s}+\hat{Z}_{s}\|_{L^{2p}(\Omega;L^{\infty}(\mathcal{O}))}^{2}$
$\qquad\qquad\qquad\qquad\qquad\qquad+\|\hat{X}_{s}+\hat{Y}_{s}\|_{L^{2p}(\Omega;L^{\infty}(\mathcal{O}))}^{2}\Big)\|\hat{Z}_{s}\|_{L^{2p}(\Omega;H)}\,\mathrm{d}s$
$+C\left(\int_{0}^{t}(t-s)^{-\frac{1}{2}}e^{-(\lambda_{1}^{2}-\epsilon_{0})(t-s)}\left(1+\|\hat{X}_{s}+\hat{Y}_{s}\|_{L^{p}(\Omega;W^{1,2}(\mathcal{O}))}^{4}+\|\hat{X}_{s}\|_{L^{p}(\Omega;W^{1,2}(\mathcal{O}))}^{4}\right)\mathrm{d}s\right)^{\frac{1}{2}}$
$\qquad\qquad\qquad\qquad\times\left\|\int_{0}^{t}e^{-\epsilon_{0}(t-s)}\left\|(-A)^{\frac{1}{2}}\hat{Y}_{s}\right\|_{L^{2}}\mathrm{d}s\right\|_{L^{p}(\Omega)}^{\frac{1}{2}}$
$\leq$ $Ce^{-\frac{\epsilon_{0}}{2}t}\left\|\hat{y}_{0}\right\|_{L^{2}}+C\sqrt{K_{3}(t)}\left(\mathbb{I}_{1}\Big(t,1/2,1/2,\big\|\hat{Z}\big\|_{L^{2p}(\Omega;L^{2}(\mathcal{O}))}^{2}\Big)\right)^{\frac{1}{2}}$
$+C\sqrt{K_{2}(t)K_{3}(t)}\left(\mathbb{I}_{1}\Big(t,0,\epsilon_{0},\big\|\hat{Z}\big\|_{L^{2p}(\Omega;L^{4}(\mathcal{O}))}^{4}\Big)\right)^{\frac{1}{4}}.$
This completes the proof. ∎
@@
Suppose that Assumptions 2.2-2.4 hold, then the transition semigroup $\{P_{t}\}_{t\in[0,\infty)}$ of SCHE (1.1) possesses the Feller property.
Proof.
For any $t\in[0,\infty)$, it follows from (3.23) and Proposition 3.4 that
$\left\|u(\hat{u}_{0};t,\cdot)-u(\tilde{u}_{0};t,\cdot)\right\|_{L^{2}(\Omega;H)}\leq Ce^{-\frac{\epsilon_{0}}{2}t}\left\|\hat{u}_{0}-\tilde{u}_{0}\right\|_{L^{2}},\quad\hat{u}_{0},\tilde{u}_{0}\in H.$ (3.29)
@@
By using the continuity and boundedness of $\phi$, (3.29), and the Hölder inequality, we obtain
@@
where ${B}^{C}$ denotes the complement of a set B{B}, and $\chi_{{B}}$ denotes its indicator function. Therefore, there exists a constant $\hat{\delta}>0$, such that, whenever $\|\hat{u}_{0}-\tilde{u}_{0}\|_{L^{2}}<\hat{\delta}$, we have $|P_{t}\phi(u_{0})-P_{t}\phi(\hat{u}_{0})|<\epsilon,$ which means $P_{t}$ is continuous. It follows from the boundedness of $\phi$ that $P_{t}$ is bounded. In summary, $\{P_{t}\}_{t\in[0,\infty)}$ possesses the Feller property ∎
3.Existence and uniqueness of invariant measure